OpenStudy (anonymous):

lex designed the model of a cylindrical water tank to be installed in a park. The model has a radius of 4 cm and a height of 12 cm. The dimensions of the actual water tanks that will be installed in the park are 5 times larger than the model. The total surface area of the actual water tanks, rounded to the nearest whole number, is _________ cm2. Use for pi. Do not type a comma in your response. For example, an answer of 1,000 would be given as 1000.

OpenStudy (saifoo.khan):

Do you know the total SA formula?

OpenStudy (saifoo.khan):

This should help, http://0.tqn.com/d/math/1/5/F/F/Cylinderr.gif

OpenStudy (anonymous):

2*pi*r ^2 + 2*pi*r*h ?

OpenStudy (saifoo.khan):

Yes. now insert the values.

OpenStudy (anonymous):

2*3.14*18 + 2*3.14*4*12 = 17338.8 is the larger one 86694?

OpenStudy (anonymous):

if you round the first cylinder to the nearest tenth

OpenStudy (saifoo.khan):

Did u found the surface area of the smaller cylinder?

OpenStudy (anonymous):

yes, the first equation is the smaller cylinder. then i multiplied it by 5

OpenStudy (saifoo.khan):

im getting 2072

OpenStudy (anonymous):

How did you get 2072?

OpenStudy (saifoo.khan):

sa of the normal cylinder = 414.8 414.8 * 5 = 2072

OpenStudy (saifoo.khan):

normal = model.

OpenStudy (anonymous):

I just got 15473 and i knoww

OpenStudy (anonymous):

lol, am I multiplying too many numbers?

OpenStudy (saifoo.khan):

might be. lol.

OpenStudy (anonymous):

What numbers did you use to get 414.8?

OpenStudy (anonymous):

Do you mind typing the equation

OpenStudy (saifoo.khan):

The same numbers you wrote, (2*3.14*18 + 2*3.14*4*12)

OpenStudy (anonymous):

lol, I got 15445.7856 and i think the 18 is supposed to be 16 because 4=r 4^2=16?

OpenStudy (saifoo.khan):

2*3.14*18 + 2*3.14*4*12 113*04 + 301.44

OpenStudy (anonymous):

OHHHHHHHHHHHHHHHHHHHHHHHHHHHHHH! I've been multiplying it!

OpenStudy (anonymous):

wait, why is it 18?

OpenStudy (saifoo.khan):

ikr.

OpenStudy (saifoo.khan):

you wrote 18. lol.

OpenStudy (saifoo.khan):

Let me do it again.

OpenStudy (saifoo.khan):

2 * 3.14 * 4^2 + 2*3.14*4*12

OpenStudy (anonymous):

Wait, I've been getting the same number. the 15445.78

OpenStudy (anonymous):

I made sure I added it

OpenStudy (saifoo.khan):

2 * 3.14 * 4^2 + 2*3.14*4*12 401.92 for model.

OpenStudy (anonymous):

2009.5?

OpenStudy (saifoo.khan):

2009.6